# Extending 6D Object Pose Estimators for Stereo Vision

### v1

@@ Abstract
Estimating the 6D pose of objects accurately, quickly, and robustly remains a difficult task. However, recent methods for directly regressing poses from RGB images using dense features have achieved state-of-the-art results. Stereo vision, which provides an additional perspective on the object, can help reduce pose ambiguity and occlusion. Moreover, stereo can directly infer the distance of an object, while mono-vision requires internalized knowledge of the object’s size. To extend the state-of-the-art in 6D object pose estimation to stereo, we created a BOP compatible stereo version of the YCB-V dataset. Our method outperforms state-of-the-art 6D pose estimation algorithms by utilizing stereo vision and can easily be adopted for other dense feature-based algorithms.
@@ 1 Introduction
In this work, we test the assumption that extending a monocular pose estimator is an effective way to greater pose accuracy and we test this hypothesis using current state-of-the-art correspondence-based algorithms. In particular our contributions include:
@@
This work is structured as follows: First, we will discuss related work. Then, we will present our dense stereo 6DOPE method and reasoning for the various fusion approaches and additional features. Next we introduce our stereo YCB-V dataset (YCB-V DS) used to train and evaluate our method. Finally, we evaluate all proposed approaches and compare our method to state-of-the-art end-to-end 6DOPE.
@@ 3 Approach
In the field of end-to-end 6DOPE, current state-of-the-art techniques are primarily focused on single images without an obvious methodology how to incorporate additional views. We propose a way to extend 6DOPE algorithms using dense feature to the stereo modality by exploring various approaches to fuse information from two views, including disparity and deep image features.
In subsequent sections, we discuss different architectures for fusing images to estimate object pose. More specifically, we evaluate the following fusion approaches: Early Fusion, Mid Fusion, Late Fusion, Double Fusion, as well as an extension with additional features. All of these architectures are depicted in Figure 1 and will subsequently be discussed in detail.
Our method builds upon the processing pipelines as found in GDRN and SO-Pose and begins with the estimator receiving detections with bounding boxes. These bounding boxes define a region of interest (RoI) and zooming into the RoI we obtain our embeddings. The next stage in the network predicts dense features such as 2D-3D correspondences, regions, and occlusion labels. Finally, a regression network determines the translation and rotation of the object from these features. Our contribution consists in exploring additional features like disparity and integrating multiple views into the existing pipeline at various locations.
@@
Feature Fusion. One of the most significant challenges in 6DOPE is pose ambiguity, which we argue, can be mitigated by incorporating additional views. To address this, we propose and evaluate four types of feature fusion of distinct views within the end-to-end network: early fusion, mid fusion, and late fusion. Additionally, we examine fusion at both the early and late stages of the network, as well as incorporating disparity features. We continue by describing our reasoning for the different approaches, and their combination.
@@ 3.1 Early Fusion
We propose merging the embeddings from the first stage in such a way that the second stage can infer 2D-3D correspondences and region classification from multiple views. To maintain the same merged channel dimension, we halve the channel dimension of the left and right embeddings. This approach allows the network to accurately identify the object in the image, particularly for objects that may appear similar from certain views but not from others. As a result, we assume improvements in region classification and 2D-3D correspondence estimation.
@@ 3.5 Early Fusion with shared Backbone Disparity Prediction
This network architecture combines various features while maintaining low complexity. To reduce computational load, we re-use our common backbone to obtain features for disparity estimation. The disparity is concatenated with the dense features in the channel dimension.
@@ 4 Evaluation
We extend state-of-the-art methods, which includes GDRN, GDRNPP, and SO-Pose, by implementing a stereo version of SO-Pose and GDRNPP, called SO-Stereo and GDRN-Stereo, respectively. For comparability with standard BOP datasets, we extend the BlenderProc [2] Pipeline to generate all necessary data and created a stereo PBR YCB-Video dataset together with real stereo recordings.
@@ 4.1 Synthetic and real Dataset
Data Generation and Processing. Our dataset includes a wide range of generated scenes with randomly sampled views, objects of interest, distraction objects, and object locations above the ground. We physically simulate the objects under gravity, which ensured a wide range of realistic occlusions in the resulting scenes. The background texture, lighting direction and intensity, viewpoints, material and reflectivity of objects were sampled randomly. By extending BlenderProc to output stereo pairs, we get perfect ground truth for the 6DOPE problem. In total, we rendered fifty thousand stereo frames consisting of RGB and depth images, including up to 15 target objects per frame. For each scene, 25 frames were obtained from random viewpoints of the scene. Annotations such as object masks and visibility percentages were computed with the BOP Toolkit. Next, the GDRN and SO-Pose repositories were used to compute additional required annotations, such as 2D-3D correspondences and occlusion correspondences. For computing 2D-3D correspondences we modified scripts from the SO-Pose repository, speeding them up using Numba, an LLVM-based Python JIT compiler, and pallelizing tasks with Python Multiprocessing. In addition, we converted the file formats from the previous papers to a compressed file format (.npz) to reduce memory requirements and save space. The SO-Pose occlusion labels were generated using scripts from the SO-Pose repository, and we again used Numba and Python Multiprocessing to accelerate the process and compress the files. Finally, we removed ground truth labels if less than 10% of the object surface was visible in either of the stereo image pairs from the dataset for stability reasons. We ended up with a training set containing 433.645 labels and a test set containing 48.080 labels. Also note that the average visibility of target objects in our dataset is only 62%, the effect of a high amount of occlusion, commonly recognized as making the 6DOPE problem more difficult.
@@ 4.2 Parameterization
Loss Function. The loss function encompasses the individual loss functions associated with GDRN, SO-Pose, and the stereo matching network. It is represented as $L=L_{Pose}+L_{CL}+L_{occ}+L_{Stereo}$ where the top-level loss functions are assigned equal weights. The parameters for the lower-level loss functions remain unaltered, consistent with the SO-Pose configuration for YCB-V.
@@
Optimization. To optimize our model, we employed the Ranger [29] optimizer with a learning rate of $0.0001$. The learning rate undergoes a warm-up phase with a factor of 0.001 during the initial 1000 iterations. In the final 27% to 72% of the training period we use learning rate annealing using the cosine function. The learning rate weights for the cross-task loss functions of SO-Pose are elevated from zero after 20% of the training, consistent with previous methodologies.
@@ 4.3 Results
In sum, all of our stereo methods outperform the monocular counterparts in the single model multi-object scenario on our heavily occluded YCB-V dataset with a significant margin.

### v2

@@ Abstract
Estimating the 6D pose of objects accurately, quickly, and robustly remains a difficult task. However, recent methods for directly regressing poses from RGB images using dense features have achieved state-of-the-art results. Stereo vision, which provides an additional perspective on the object, can help reduce pose ambiguity and occlusion. Moreover, stereo can directly infer the distance of an object, while mono-vision requires internalized knowledge of the object’s size. To extend the state-of-the-art in 6D object pose estimation to stereo, we created a BOP compatible stereo version of the YCB-V dataset called YCB-V DS and extended multiple state-of-the-art pose estimator algorithms to be applicable to stereo. Our method outperforms state-of-the-art 6D pose estimation algorithms by utilizing stereo vision and can easily be adopted for other dense feature-based algorithms.
@@ 1 Introduction
In this work, we test the assumption that extending a monocular pose estimator to stereo vision is an effective way to greater pose accuracy and we test this hypothesis using current state-of-the-art correspondence-based algorithms. In particular our contributions include:
@@
This work is structured as follows: First, we will discuss related work. Then, we will present our dense stereo 6DOPE method and reasoning for its design, motivating the applicability to other, similar algorithms. Next we introduce our stereo YCB-V dataset (YCB-V DS) used to train and evaluate our method. Finally, we evaluate all proposed approaches and compare our method to state-of-the-art end-to-end 6DOPE.
@@ 3 Approach
In the field of end-to-end 6DOPE, current state-of-the-art techniques are primarily focused on single images without an obvious methodology how to incorporate additional views. We propose a way to extend 6DOPE algorithms using dense features to the stereo modality by exploring various approaches to fuse information from two views and introducing disparity features.
In subsequent sections, we discuss different architectures for fusing images to estimate object pose. Fusion is the process of combining the signals of both images at different stages of the processing pipeline. More specifically, we evaluate the following fusion approaches: Early Fusion, Mid Fusion, Late Fusion, Double Fusion, as well as an extension with additional features. All of these architectures are depicted in Figure 1 and will subsequently be discussed in detail.
Our method builds upon the processing pipelines as found in GDRN and SO-Pose and begins with the estimator receiving detections with bounding boxes. These bounding boxes define a region of interest (RoI) and zooming into the RoI we obtain our embeddings. The next stage in the network predicts dense features such as 2D-3D correspondences, regions, and occlusion labels. Finally, a regression network determines the translation and rotation of the object from these features. Our contribution consists in exploring additional features like disparity and integrating multiple views into the existing pipeline at various stages.
@@
Feature Fusion. One of the most significant challenges in 6DOPE is pose ambiguity, which we argue, can be mitigated by incorporating additional views. To address this, we propose and evaluate the influence of fusing the extracted features of two images at different stages of the pipeline. We continue by describing our reasoning for the different approaches, and their combination.
@@ 3.1 Early Fusion
We propose merging the embeddings from the first stage in such a way that the second stage can infer 2D-3D correspondences and region classification from multiple views. To this end the feature maps from the left and the right view get concatenated after the backbone but before the decoder as shown in Figure 1 a). To maintain the same merged channel dimension, we halve the channel dimension of the left and right embeddings. This approach should allow the network to accurately identify the object in the image, particularly for objects that may appear similar from certain views but not from others. As a result, we assume improvements in region classification and 2D-3D correspondence estimation.
@@ 4 Evaluation
We extend state-of-the-art methods, which includes GDRN, GDRNPP, and SO-Pose, by implementing a stereo version of SO-Pose and GDRN, called SO-Stereo and GDRN-Stereo, respectively. For comparability with standard BOP datasets, we extend the BlenderProc [2] Pipeline to generate all necessary data and created a stereo PBR YCB-Video dataset called YCB-V DS together with real stereo recordings.
@@ 4.1 Synthetic and real Dataset
Data Generation and Processing. Our dataset includes a wide range of generated scenes with randomly sampled views, objects of interest, distraction objects, and object locations. We physically simulate the objects under gravity, which ensures a wide range of realistic occlusions in the resulting scenes. The background texture, lighting direction and intensity, viewpoints, material and reflectivity of objects were sampled randomly. By extending BlenderProc to output stereo pairs, we get perfect ground truth for the 6DOPE problem. In total, we rendered fifty thousand stereo frames consisting of RGB and depth images, including up to 15 target objects per frame. For each scene, 25 frames were obtained from random viewpoints of the scene. Annotations such as object masks and visibility percentages were computed with the BOP Toolkit. Next, the GDRN and SO-Pose repositories were used to compute additional required annotations, such as 2D-3D correspondences and occlusion correspondences. For computing 2D-3D correspondences we modified scripts from the SO-Pose repository, speeding them up using Numba, an LLVM-based Python JIT compiler, and pallelizing tasks with Python Multiprocessing. In addition, we converted the file formats from the previous papers to a compressed file format (.npz) to reduce memory requirements and save space. The SO-Pose occlusion labels were generated using scripts from the SO-Pose repository, and we again used Numba and Python Multiprocessing to accelerate the process and compress the files. Finally, we removed ground truth labels if less than 10% of the object surface was visible in either of the stereo image pairs from the dataset for stability reasons. We ended up with a training set containing 433.645 labels and a test set containing 48.080 labels. Also note that the average visibility of target objects in our dataset is only 62%. High amount of occlusion is commonly recognized as increasing the difficulty of the 6DOPE problem, making our dataset quite challenging.
@@ 4.2 Parameterization
Loss Function. The loss function encompasses the individual loss functions associated with GDRN, SO-Pose, and the stereo matching network (if used). We do not change the weighting of individual loss terms compared to the baseline algorithms.
@@
Optimization. To optimize our model, we employed the Ranger [29] optimizer with a learning rate of $0.0001$. The learning rate undergoes a warm-up phase with a factor of 0.001 during the initial 1000 iterations. In the final 27% to 72% of the training period we use learning rate annealing using the cosine function.
@@ 4.3 Results
In sum, all of our stereo methods outperform the monocular counterparts in the single model multi-object scenario on our heavily occluded YCB-V DS dataset with a significant margin.